MC Qu. 110 Excerpts from Dowling Company's... Excerpts from Dowling Company's December 31, 2018 and 2017, financial statements and key ratios are presented below (all numbers are in millions): Dowling's 2018 average collection period is?

2018 2017
Accounts receivable (net) $ 22 $ 14
Net sales $ 126 $ 111
Cost of goods sold $ 71 $ 66
Net income $ 22 $ 17
Inventory turnover 5.75
Return on assets 11.30 %
Equity Multiple 2.47
Answer #1

Average accounts receivable = (Beginning accounts receivable + Ending accounts receivable)/2

= (14+22)/2

= 18

Accounts receivable turnover = Net credit sales/Average accounts receivables

= 126/18

= $7 million

Average collection period = 365/Accounts receivable turnover

= 365/7

= 52 days

Kindly comment if you need further assistance. Thanks‼!